# Shuofei Qiao♠♡, Pengjun Xie, Fei Huang, Huajun Chen, Ningyu Zhang

### v2

@@ Introduction
  • Validation: After every $t$ tasks, only the trajectories of successfully completed tasks are retained and converted into procedural memories for storage.
@@
As depicted in Figure 3, we systematically divided the tasks within our testbed into several distinct groups, with each group comprising a diverse set of individual tasks. Upon the completion of tasks within each group, we employed the previously described strategies to construct, store, and update the procedural memory. The experimental results reveal a clear trend: as we sequentially progress through more groups and iteratively refresh the memory, all strategies contribute to improved performance on subsequent tasks. Specifically, this is reflected not only in higher overall scores but also in a reduction in the number of steps required to complete the tasks.
@@
Analysis
Procedural Memory Boosts Accuracy and Cuts Trials.
Figure 5 presents a case study demonstrating how Procedural Memory enhances both accuracy and efficiency. In the absence of Procedural Memory, facing a complex task that has not been performed before, there are usually two situations. In the first scenario (left), the model repeatedly attempts illegal or incorrect actions, causing the context to become increasingly complex and eventually exceeding the model’s understanding capacity. In the second scenario (middle), after multiple attempts, the model completes the task but at a cost significantly higher than the optimal path.
@@
For a procedural memory constructed from a strong model in an offline memory library, we aim to verify whether this form of procedural memory can be effectively transferred to other models, or even weaker models. This exploration underscores the significance of memory transfer, as it could potentially enhance the adaptability and efficiency of various models by leveraging the knowledge and experience encapsulated within the strong model’s memory structure. As shown in Figure 4 (b), procedural memory generated by GPT-4o was employed by Qwen2.5-14B. On the Travel Plan benchmark, the 14 billion parameter model raised its task completion rate by 5% and cut the average number of steps by 1.6. Similar gains, both in success rate and trajectory length, appeared on ALFWorld. These outcomes confirm that procedural knowledge from a stronger model can be distilled into a reusable memory bank and transferred to a smaller model with minimal overhead, giving that smaller model a clear boost in task solving ability. Moreover, by leveraging procedural memory transfer, we can rapidly migrate the experiential knowledge that one model has acquired to another, which is highly beneficial for agents as they adapt to new tasks with greater efficiency and robustness.
Scaling Memory Retrieval Improves Agent Performance.
While our main experiment has already demonstrated that procedural memory improves an agent’s task accuracy and reduces the number of steps required, vector-based storage and retrieval confer an advantage over human procedural memory: they can be scaled both in total capacity and in the number of memories retrieved. To investigate whether an agent’s performance continues to rise as the procedural-memory store and the number of retrieved memories increase, we designed a set of follow-up experiments. As showned in Figure 4 (b), as the number of retrieved procedural memories increases, the agent’s performance also improves steadily, exhibiting an upward trend followed by a plateau. However, retrieving too many memories can lead to a decline in the agent’s performance. This is because excessive retrieval can affect the context length and also introduce less accurate procedural memories, which can interfere with the overall effectiveness.
Conclusion and Future Work
We introduce $Mem^{p}$, a task-agnostic framework that elevates procedural memory to a core optimization target in LLM-based agents. By systematically studying strategies for memory construction, retrieval, and updating, $Mem^{p}$ enables agents to distill, reuse, and refine their own past experiences across diverse, long-horizon tasks. Empirical results on housework automation and information-seeking benchmarks show that leveraging procedural memory significantly boosts task success rates and efficiency. Beyond improving individual episodes, $Mem^{p}$ supports continual learning and robust generalization, marking a step toward self-improving, resilient agents.
In our experiments, $Mem^{p}$ has achieved promising results in both construction and retrieval. Moving forward, we plan to enhance this work in several ways. Firstly, we will develop more diverse retrieval strategies. The current approach involves constructing different keys for vector-based retrieval. However, traditional methods like BM25 could also be explored to retrieve precise memories more effectively. Secondly, in $Mem^{p}$, we currently rely on the standard reward signals provided by the benchmark. However, in real-world scenarios, many tasks do not have clear reward signals, making it difficult for the agent to determine whether a task has been completed successfully. In such cases, using a large language model (LLM) as a judge to assess task completion could be a viable solution. This would transform the agent’s lifecycle into a continuous loop of executing tasks, self-assessing completion, building memories, and then proceeding to new tasks.

### v3

@@ 4.3 Memory Update
    Validation: After every $t$ tasks, the system performs a selective consolidation step: only the trajectories that terminated in successful task completion are extracted, abstracted into compact, symbolic procedural memories. Trajectories that ended in failure, along with any redundant or noisy data, are discarded.
@@
As depicted in FigureĀ 3, we systematically divided the tasks within our testbed into several distinct groups, with each group comprising a diverse set of individual tasks. Upon the completion of tasks within each group, we employed the previously described strategies to construct, store, and update the procedural memory. The experimental results reveal a clear trend: as we sequentially progress through more groups and iteratively refresh the memory, all strategies contribute to improved performance on subsequent tasks.
@@
4.4 Comparison
Figure 4: Performance of several baselines using GPT-4o on ALFWorld
FigureĀ 4 compares the performance and efficiency of different baselines using GPT-4o on ALFWorld. React performs the worst, with low success rates and the highest number of steps, indicating inefficient interaction. Expel and AWM achieve progressively better performance, with AWM outperforming Expel on both splits. $Mem^{p}$ delivers the best results, attaining the highest dev and test success rates while requiring the fewest steps. This shows that memory-augmented mechanisms significantly improve both task effectiveness and execution efficiency in long-horizon embodied agent tasks.
5 Analysis
Procedural Memory Boosts Accuracy and Cuts Trials.
FigureĀ 6 presents a case study demonstrating how Procedural Memory enhances both accuracy and efficiency. In the absence of Procedural Memory, facing a complex task that has not been performed before, there are usually two situations. In the first scenario, the model repeatedly attempts illegal or incorrect actions, causing the context to become increasingly complex and eventually exceeding the model’s understanding capacity. In the second scenario, after multiple attempts, the model completes the task but at a cost significantly higher than the optimal path.
@@ 5 Analysis
For a procedural memory constructed from a strong model in an offline memory library, we aim to verify whether this form of procedural memory can be effectively transferred to other models, or even weaker models. This exploration underscores the significance of memory transfer, as it could potentially enhance the adaptability and efficiency of various models by leveraging the knowledge and experience encapsulated within the strong model’s memory structure. As shown in FigureĀ 5 (b), procedural memory generated by GPT-4o was employed by Qwen2.5-14B. On the Travel Plan benchmark, the 14 billion parameter model raised its task completion rate by 5% and cut the average number of steps by 1.6. Similar gains, both in success rate and trajectory length, appeared on ALFWorld. These outcomes confirm that procedural knowledge from a stronger model can be distilled into a reusable memory bank and transferred to a smaller model with minimal overhead, giving that smaller model a clear boost in task solving ability.
Scaling Memory Retrieval Improves Agent Performance.
While our main experiment has already demonstrated that procedural memory improves an agent’s task accuracy and reduces the number of steps required, vector-based storage and retrieval confer an advantage over human procedural memory: they can be scaled both in total capacity and in the number of memories retrieved. To investigate whether an agent’s performance continues to rise as the procedural-memory store and the number of retrieved memories increase, we designed a set of follow-up experiments. As shown in FigureĀ 5 (b), as the number of retrieved procedural memories increases, the agent’s performance also improves steadily, exhibiting an upward trend followed by a plateau. However, retrieving too many memories can lead to a decline in the agent’s performance. This is because excessive retrieval can affect the context length and also introduce less accurate procedural memories, which can interfere with the overall effectiveness.
6 Conclusion and Future Work
We introduce $Mem^{p}$, a task-agnostic framework that elevates procedural memory to a core optimization target in LLM-based agents. By systematically studying strategies for memory construction, retrieval, and updating, $Mem^{p}$ enables agents to distill, reuse, and refine their own past experiences across diverse, long-horizon tasks and supports continual learning and robust generalization, marking a step toward self-improving, resilient agents.
In our experiments, $Mem^{p}$ has achieved promising results in both construction and retrieval. Moving forward, we plan to enhance this work in several ways. Firstly, we will develop more diverse retrieval strategies. The current approach involves constructing different keys for vector-based retrieval. Secondly, in $Mem^{p}$, we currently rely on the standard reward signals provided by the benchmark. In such cases, using a LLM as a judge to assess task completion could be a viable solution. This would transform the agent’s lifecycle into a continuous loop of executing tasks, self-assessing completion, building memories, and then progress.
Limitations
$Mem^{p}$ is still constrained by two key limitations: 1. retrieval is restricted to vector-similarity search with manually crafted keys—other classic, potentially more precise methods like BM25 have not been incorporated; 2. the framework depends on explicit, benchmark supplied reward signals, so it cannot yet judge task success in real world settings where such rewards are sparse or absent.
Appendix
Appendix A Use of AI Assistant
We affirm that Large Language Models are employed solely as an assisted tool to refine wording and sentence structure during our paper writing process. Their use in the experiments is strictly for scientific research purposes, and all such usage has been explicitly documented in our Experimental Settings and Reproducibility Statement. No other reliance on LLMs is involved in this work.
Appendix B Case Study
Detailed case studies are provided in FigureĀ 6.
@@ Appendix B Case Study
Appendix C Evaluation Details